OpenStudy (anonymous):

If sinx=4/5 and x is in the first quadrant, how would I find each of the following?

OpenStudy (anonymous):

cos2x, sin2x, tan2x step by step i'm kind of lost

OpenStudy (tkhunny):

By formula. If x is in the 1st Quadrant, then 2x is in the 1st or 2nd Quadrant. Thus, cosine can be positive or negative, but sine will always be positive. \(cos(2x) = 1 - 2\sin^{2}(x)\) \(\cos(x) = \pm\sqrt{1 - \sin^{2}(x)}\) Consider other formulas that might be useful.

OpenStudy (anonymous):

So where would sinx=4/5 come in?

OpenStudy (tkhunny):

Both the formulas I have written. There may be other places as you think about the rest of the problem. For one thing, you are going to have to decide which sine is correct on that \(\cos(2x)\) formula.

OpenStudy (anonymous):

Okay, thank you! :)